OpenStudy (anonymous):

MEDALSSS!! Solve x3 = 1 over 8.

OpenStudy (misty1212):

\[x^3=\frac{1}{8}\]?

OpenStudy (anonymous):

yes @misty1212

OpenStudy (misty1212):

do it top and bottom the cubed root of one is one the cubed root of eight is two (since 2^3=8\)

OpenStudy (anonymous):

1/4?

OpenStudy (misty1212):

no dear \(\frac{1}{2}\)

OpenStudy (anonymous):

so its 4/8 ?